OpenStudy (anonymous):

A culture started with 5,000 bacteria. After 4hours, it grew to 6,000 bacteria. Predict how many bacteria will be present after 17ours. Round your answer to the nearest whole number. P=Ae^kt

OpenStudy (anonymous):

I think I got the answer, I want to know if it's right.

OpenStudy (anonymous):

easy way or hard way

OpenStudy (anonymous):

Easy way

OpenStudy (sriram):

e is approx 2.7 right??

OpenStudy (anonymous):

A=5000*(1.1)^17/6 =6550 ?

OpenStudy (anonymous):

Is it wrong?

OpenStudy (anonymous):

i will do it the easy way. \[6000\div 5000=1.2\] so it get 20% in 4hours. formula is \[5000(1.2)^{\frac{t}{4}}\] i replace t by 17 to get \[5000(1.2)^{\frac{17}{4}}=10851\]

OpenStudy (anonymous):

Is it wrong?

OpenStudy (anonymous):

Im still confused :(

OpenStudy (anonymous):

@ilovemath your denominator should be 4, not 6

OpenStudy (anonymous):

and it increased by 20% not 10%, so you should use 1.2 not 1.1

OpenStudy (anonymous):

first line is A culture started with 5,000 bacteria. After 4hours, it grew to 6,000 bacteria. so we know that 6000/5000 = 1.2 and we know it increases by this factor every 4 hours, not every six hours. so you should have \[5000(1.2)^{\frac{t}{4}}\] for your fomula

OpenStudy (anonymous):

unconfused or still stuck. if stuck let me know

OpenStudy (anonymous):

Is it wrong?

OpenStudy (anonymous):

I think I get it, im going to try & do one & im going to to post it up, to see if i got it right (:

OpenStudy (anonymous):

ok but did you understand the formula i wrote? because that is the right one. it is not \[5000(1.1)^{\frac{17}{6}}\] but should be \[5000(1.2)^{\frac{17}{4}}\]

OpenStudy (anonymous):

so what does it come out to?